# Beehive: A Flexible Network Stack for Direct-Attached Accelerators

### v3

@@ 5.2. Consensus Witness
Consensus algorithms are an essential part of many deployed distributed systems as they enable a strictly consistent order for stateful client operations even in the face of failures and message delays/retransmissions. Most consensus algorithms (liskov12vr, ; paxoslive, ; raft, ) follow a common pattern: an elected leader proposes an order for arriving client requests, verifies with a set of replicas that it is still leader, and commits the request. It then performs any necessary application logic (e.g., to update state), replies back to the client, and informs the other replicas, so that they can also perform the application logic in the same order.
@@
A common type of application built on top of consensus is a key-value (KV) store. To achieve higher throughput, the key space is often sharded with a leader and replica set for each slice. However, even with sharding, consistent reads can be expensive, because the leader must validate, each time, that it is still the leader before replying with the value stored with the key. As a result, it is common in practice to configure the system to return stale reads, allowing the leader to reply immediately (zookeeper, ; crooksweak, ). This places a burden on the client developer to handle the (rare) case where a failover can lead to inconsistent client data.
@@
In our evaluation, we show that a consensus accelerator can help reduce the cost of consistency. Our accelerator operates as a witness, that is, it only validates the leader and tracks the operation order; it does not execute client operations. Single node fault tolerance can be achieved with one leader, one witness, and one replica. To add further fault tolerance, we add additional witnesses and replicas. For example, two-node fault tolerance can be achieved with one leader, two witnesses, and two replicas. To validate a read or write operation, the leader only needs to receive a verification from the witnesses before replying to the client. The witness can be designed in hardware to reply with low and reliable latency.
@@
Our witness protocol is based on a modified version of the Viewstamped Replication (VR) used in previous studies of high-performance consensus (specpaxos, ). VR witnesses are integrated into Beehive as UDP applications. To handle multiple shards, we use one VR witness tile per shard. Unlike the RS encoder, the VR witness is not stateless and requests for a shard must always go to the same tile. We distribute work to the VR tiles by matching on the destination port number.
5.3. TCP live migration
Live TCP migration is useful for load balancing and server consolidation (prism_nsdi, ). It allows TCP connections to remain active while virtual machines are migrated in response to offered load, such as during scale down where connections are consolidated onto fewer nodes. To support this, we integrate a network address translation (NAT) service into Beehive, as described in Section 4.5. Specifically, we integrate the NAT tiles between the TCP and IP tiles both on the receive and transmit paths to perform the translation. We also add a new tile to the control plane to handle the management interface. Then, we set up the routing tables of the IP tiles on the receive path to send to the NAT tile, and the NAT tile to send to the TCP tile. We also modify the routing tables of the TCP tiles on the transmit tile to send packets to the NAT tile instead of the IP tile. The internal controller is integrated as an application with the TCP stack.
We can insert this new functionality without any modifications to either the IP or TCP protocol implementations, because the protocols are separated into their own tiles.
Figure 6. Packet size vs. goodput for a UDP echo application. Beehive with one instance of every protocol tile provides higher bandwidth than Demikernel at all packet sizes.
6. Evaluation
@@
Our evaluation tests Beehive’s ability to support scalability, low latency, and flexibility in a range of network stack configurations. We begin by evaluating Beehive with UDP and TCP microbenchmarks designed to test RPC performance and then evaluate three different application use cases: Reed-Solomon encoding acceleration, and Viewstamped Replication acceleration, and TCP connection live migration.
Figure 7. Packet size vs. goodput for Beehive and Linux TCP send. The (CPU send/FPGA receive) is omitted, as it is approximately the same as (CPU send/CPU receive) due to the CPU send path being the bottleneck.
6.1. Flexibility
As a quantitative proxy for flexibility, we count the lines of code (LoC) required to insert an additional instance of an implemented service (network function or application) into the design for our three designs. Results are shown in Table 1. We count both NAT tiles for the TCP migration service.
Table 1. Lines of code per new tile Beehive  for end-to-end applications. XML configuration numbers are given as LoC for declaring the tile plus the LoC to add it as a destination.
Lines of Code
XML Configuration Verilog Top Level
Reed-Solomon 25 + 6 13
Viewstamped Replication 18 + (6 $\times$ # of UDP tiles) 17
TCP Migration 2 $\times$ (34 + 6) 2 $\times$ 15
6.2. Setup
@@
We use Vivado 2021.2 for building our FPGA images. Beehive is configured on an Alveo U200 at 250 MHz. The FPGA and the clients are connected to an Edgecore Arista DCS-7060CX-32S-R 100G switch with jumbo frames enabled. We use five machines during evaluation with Turboboost disabled. All of them have Mellanox ConnectX-5 100G NICs and are running Ubuntu 20.04. Two machines have Intel Xeon Gold 6226R CPUs; the other three machines have Intel Xeon Gold 5218 CPUs.
@@
In experiments where energy is measured, we use the RAPL counters on the CPUs and the Alveo CMS registers on the FPGA. For CPU energy experiments we use a two-socket machine, so we run all the application and network processing code on one socket and poll the counters from the other socket. We only use RAPL’s CPU counters, which is an underestimate as we do not include DRAM energy or network interface energy. On the FPGA, we use the Corundum framework to read the CMS registers that report instantaneous power and current usage (alveo_cms_register_space, ). We poll these counters every second to calculate energy over the benchmarking period.
6.3. UDP Echo
@@
Throughput: We first compare Beehive’s UDP echo goodput to Demikernel, an optimized DPDK network stack, on different packet sizes. Three Intel Gold 5128 machines act as clients; for the CPU experiments, the server runs on the Intel Gold 6226R machine. To stay comparable, we configure one echo application tile on the FPGA and one CPU core in Demikernel. While the CPU server uses DPDK, clients use the standard Linux network stack. We spawn the number of threads that yields highest server bandwidth for that packet size; threads send in an open-loop manner. The results of this experiment are shown in Figure 6.
@@
Beehive achieves line rate with 1024-byte packets while the optimized CPU stack remains far below maximum bandwidth even with jumbo frames. The difference in performance is especially pronounced at small packet sizes where Beehive is able to sustain echoing 9 Gbps of 64-byte packets (18392 KReq/s) whereas single core Demikernel provides 0.3 Gbps (584 KReq/s), a 31$\times$ speedup.
@@
Latency: For our latency experiment, we use Beehive and a single client thread to ping-pong a single 1-byte UDP packet. We record the latency by tagging the packet with a timestamp when it enters the network stack at the Ethernet parsing layer, taking another timestamp when it finally exits the Ethernet layer on transmit, and recording both timestamps into a log which we read back over the network. The latency through Beehive is 368 ns (92 cycles).
@@
Table 2. Energy consumption and goodput for Reed-Solomon encoding using Beehive versus CPU for 1, 2, 3 and 4 application instances.
@@
Energy efficiency 22x 20x 20x 16x
@@
Speedup 7.5x 7.8x 7.5x 7.8x
6.4. TCP throughput
To characterize the throughput performance of our TCP engine, we run a single-connection experiment and measure unidirectional sending and receiving performance across a range of packet payload sizes. Because Demikernel’s TCP implementation is optimized for latency, it performs worse than Linux on this experiment, so we configure Demikernel to use Linux TCP as its backend. The sending application sits in a tight loop, submitting data into the network stack as fast as possible; the receiver pulls data out of the network stack without doing further processing on it.
@@ 6.4. TCP throughput
6.5. Reed-Solomon Encoding Acceleration
To evaluate Beehive’s scaling architecture, we evaluate a duplicated Reed-Solomon (RS) encoding accelerator on Beehive versus a CPU implementation of the same algorithm. The client sends blocks of 4 KB to the encoder using UDP; the accelerator replies with 1K of erasure data. This could be organized into an (8,2) stripe for double fault tolerance. One instance of the Reed-Solomon encoder can consume data at 15 Gbps; our FPGA has room for four encoder instances, which consume data at 62 Gbps as shown in Table LABEL:tab:rs_encode_scaling. For comparison, we use the open-source Reed-Solomon encoding implementation from BackBlaze (backblaze_rs_encode, ) running on CPUs which we then duplicate across cores.
@@ 6.5. Reed-Solomon Encoding Acceleration
6.6. Viewstamped Replication Witness Acceleration
@@
Figure 8. Experimental setups for VR evaluation
We next turn to a latency-sensitive application, evaluating Beehive hosting a viewstamped replication (VR) witness appliance. We first evaluate the witness on a single shard. We then take advantage of Beehive’s ability to duplicate both internal components and applications to host a 4-shard witness appliance. We also duplicate protocol tiles to prevent them from becoming a bottleneck
For all experiments, we evaluate a three-node VR configuration as shown in Figure 8 with either the FPGA or CPU serving as a witness. The other nodes run on CPU. The CPU VR replicas run on Intel Xeon Gold 5218 CPUs. Client threads run on Intel Xeon Gold 6226R CPUs. Each client thread sends one request and waits for a reply before sending another. We use UDP as our transport protocol, because VR does not assume reliable message delivery.
We evaluate our VR accelerator with a replicated key-value store application with 64-byte keys and 64-byte values. We use a read-write mix of 90% reads and 10% writes and a uniform key distribution. We set the IRQ affinity to the CPU socket where the replicas are running.
We test both single-shard and multishard configurations. Latency is measured at the clients as the time between the initial request and the eventual response. The shard leaders are distributed evenly between two CPU machines. The CPU witness(es) run on a separate server to allow us to measure the energy used by a CPU witness appliance.
Latency & throughput: We plot latency versus throughput for differing numbers of shards in Figure 9. We increase offered load by increasing the number of client threads sending requests to the leader. The results are shown in Figure 9. The system using the FPGA witness can provide up to 1.14$\times$ more per-core throughput and up to 1.13$\times$ lower median latency.
Figure 9. Latency vs. throughput for the key-value store workload varying the number of shards and client threads. The FPGA witness consistently outperforms the equivalent CPU cores in both latency and throughput.
Table 3. Energy per operation (measured at the witness) and performance metrics (measured at the clients) at the circled points in Figure 9
@@
Energy: For each shard, we take the median energy measurement, throughput, median latency, and 99th-percentile (p99) latency at each circled point in Figure 9. The results are shown in Table 3. The FPGA is between 2.07$\times$ and 2.32$\times$ more energy efficient per operation compared to the CPU while providing better overall throughput and latency to key-value store clients.
6.7. TCP Connection Live Migration
To demonstrate Beehive’s flexibility, we add support for live migration of a remote TCP endpoint to our hardware TCP stack by adding tiles and reusing other components.
Figure 10. Request throughput over time. A single client sends requests every 100 $\mu$s to the server. Migration occurs around $t=0.07$ s. From $t=0$ s until migration, the TCP connection is on the first client. Throughput dips during migration, then recovers after 500 $\mu$s when the connection is brought up on the second client.
For the remote TCP endpoint, we use Demikernel (demikernel_sosp, ), because their TCP implementation is able to pause a connection, serialize the current state, and reinstall the state in the Demikernel network stack on a different node. Our test application is an echo RPC server. During the experiment, we run one closed-loop client on a CPU generating 64-byte payloads every 100 $\mu$s. The throughput around the migration event is shown in Figure 10. The migration latency, measured on the FPGA as the time between the last request from the first client and the first request from the second client, is 500 $\mu$s.
6.8. Hardware Resource Utilization
@@
Table 4. FPGA resource utilization of selected modules.
LUTs BRAM
Beehive UDP full 57272 40
UDP RX Tile 9780 9.5
         Router 5961 0
         NoC Message Processing 515 0
         UDP RX Processing 2984 9.5
TCP RX Engine 11672 16.5
@@
The hardware utilization of the Beehive infrastructure is shown in Table 4. For the UDP stack used in Section 6.3, Beehive components use 4% of the LUTs available on the Alveo U200 and 2% of the BRAMs. In a tile, a router uses around 6000 LUTs, twice the size of the UDP processing. For comparison with a more complex module, we include the utilization of the TCP receive path.

### v5

@@ VI-B Consensus witness
To demonstrate how Beehive performs in a latency-sensitive, communication-intensive application, we construct a consensus system that uses FPGA-accelerated witness nodes. Consensus algorithms are an essential part of many deployed distributed systems as they enable a strictly consistent order for stateful client operations even in the face of failures and message delays/retransmissions. Most consensus algorithms [51, 14, 60] follow a common pattern: an elected leader proposes an order for arriving client requests, verifies with a set of replicas that it is still leader, and commits the request. It then performs any necessary application logic (e.g., to update state), replies back to the client, and informs the other replicas, so that they can also perform the application logic in the same order. Because there are multiple round-trips between nodes to complete one round of consensus, message-handling latency and tail latency are especially important [82].
@@
A common type of application built on top of consensus is a key-value (KV) store. To achieve higher throughput, the key space is often sharded with a leader and replica set for each slice. However, even with sharding, consistent reads can be expensive, because the leader must validate, each time, that it is still the leader before replying with the value stored with the key. As a result, it is common in practice to configure the system to return stale reads, allowing the leader to reply immediately [35, 20, 37]. This places a burden on the client developer to handle the (rare) case where a failover can lead to inconsistent client data.
@@
In our evaluation, we show that a consensus accelerator can help reduce the cost of consistency [38], especially in a multi-shard setting. Our accelerator operates as a witness, that is, it only validates the leader and tracks the operation order; it does not execute client operations. Single node fault tolerance can be achieved with one leader, one witness, and one replica. To add further fault tolerance, we add additional witnesses and replicas. For example, two-node fault tolerance can be achieved with one leader, two witnesses, and two replicas. To validate a read or write operation, the leader only needs to receive a verification from the witnesses before replying to the client. The witness can be designed in hardware to reply with low and reliable latency.
Prior work [38, 37] has demonstrated full offload of consensus and application logic to an FPGA. We target a use case where application logic remains on CPUs and only a portion of the consensus protocol is run on Beehive. Importantly, this requires no change to the CPU-based application running on top of the consensus engine. This is advantageous as consensus algorithms are commonly used as a building block in larger distributed systems, so this allows accelerated consensus to be used without requiring the whole application to be ported to hardware. We also demonstrate how Beehive can be used to scale a consensus system to support multiple shards, which previous work did not explore.
@@
Our witness protocol is based on a modified version of the Viewstamped Replication (VR) used in previous studies of high-performance consensus [63]. VR witnesses are integrated into Beehive as UDP applications. To handle multiple shards, we use one VR witness tile per shard. Unlike the RS encoder, the VR witness is not stateless and requests for a shard must always go to the same tile. We distribute work to the VR tiles by matching on the destination port number.
VII Evaluation
@@
Our evaluation tests Beehive’s ability to support scalability, low latency, and flexibility in a range of network stack configurations. We begin by evaluating Beehive with UDP and TCP microbenchmarks designed to test RPC performance and then evaluate two case studies: Reed-Solomon encoding acceleration and Viewstamped Replication acceleration.
VII-A Setup
@@
We use Vivado 2021.2 for building our FPGA images. Beehive is configured on an Alveo U200 at 250 MHz. The FPGA and the clients are connected to an Arista DCS-7060CX-32S-R 100G switch with jumbo frames enabled. We use five machines during evaluation with Turboboost disabled. All of them have Mellanox ConnectX-5 100G NICs and are running Ubuntu 20.04. Two machines have Intel Xeon Gold 6226R CPUs; the other three machines have Intel Xeon Gold 5218 CPUs.
@@
In experiments where energy is measured, we use the RAPL counters on the CPUs and the Alveo CMS registers on the FPGA. For CPU energy experiments we use a two-socket machine, so we run all the application and network processing code on one socket and poll the counters from the other socket. We only use RAPL’s CPU counters, which is an underestimate as we do not include DRAM energy or network interface energy. On the FPGA, we use the Corundum framework to read the CMS registers that report instantaneous power and current usage [76]. We poll these counters every second to calculate energy over the benchmarking period.
VII-B Baselines
Hardware Network Stacks (PANIC and Limago): We compare against PANIC, an FPGA-based smartNIC framework, for our UDP echo microbenchmark. We are unable to compare against PANIC for our other applications using UDP, because they require scaling to more tiles than PANIC supports, and PANIC’s memory allocation makes it unwieldy to generate responses of a different size than the request. We also cannot compare against PANIC for our TCP microbenchmark, because it cannot support reliable transport applications. We also evaluate in PANIC’s original simulation evaluation infrastructure, because their released code does not include an FPGA flow. We integrated it into Corundum as suggested in the documentation, but we were unable to get it to meet timing for the Alveo U200. While they used an ADM-PCIE-9V3 [24], both our board and theirs have 16nm FPGA parts. The Alveo’s FPGA part is also comparable in resources available to the ADM-PCIE-9V3. For these reasons, we think the comparison is fair.
We compare against Limago, an HLS TCP stack, for our hardware utilization. We were unable to run benchmarks on it, because the QSFP links did not come up when the image was put on the board.
Software Network Stacks (Demikernel and Linux): We also compare against Demikernel [79], an optimized DPDK network stack, in cases where it is faster than Linux. This is only the case in the UDP echo benchmark. Otherwise, we compare against Linux’s network stack.
Figure 7: Packet size vs. goodput for a UDP echo application. Beehive and CALM perform almost identically across all packet sizes and outperform Demikernel.
VII-C UDP echo
(a) Setup for Beehive’s UDP stack
(b) Setup for the fixed pipeline UDP stack
Figure 8: UDP stacks for the echo microbenchmark
Throughput: We compare UDP echo goodput for Beehive (shown in Figure 8(a)) and Demikernel on different packet sizes. We also evaluate these against an FPGA with a pipelined UDP network stack design where the protocol engines are connected directly (shown in Figure 8(b)), and a UDP network stack implemented within the PANIC framework which we will refer to as CALM.
@@ VII-C UDP echo
In our experiments, the Demikernel server runs on an Intel Gold 6226R machine, and we use three Intel Gold 5128 machines as clients using the standard Linux network stack. We spawn the number of client threads that yields the highest server bandwidth for that packet size, and they send in an open-loop manner. We give the server a single core to compare against Beehive’s single application tile.
For Beehive, we run a packet generator on another U200 FPGA. This is because the client machines used for the CPU experiments cannot generate enough traffic to saturate the FPGA. We use 7 tiles in total: we separate the Ethernet, IP, and UDP layers, and then we separate their receive and transmit paths for 6 tiles and then one tile for the application.
For CALM, we implement a UDP echo server within its framework starting from their publicly available code [59]. We use 3 tiles to implement the echo server: one providing a fixed UDP receive path, one providing the application, and one providing a fixed UDP send path. We were unable to modify PANIC to support more than 8 tiles, only 4 of which are available for user functionality, so we could not make every layer into a tile as we do in Beehive. We note that this means it is less flexible than Beehive’s network stack, because we lose the opportunity to easily insert network functions or alternate protocols alongside the UDP paths.
@@
Figure 7 shows our throughput benchmark results. Beehive and CALM provide similar performance despite Beehive having more tiles. Both achieve line rate at 1024 byte packets. Beehive on FPGA levels out at this point, because the actual Ethernet link has a maximum bandwidth of 100 Gbps. However, in simulation, both Beehive and CALM continue to scale to the theoretical maximum of 128 Gbps. The pipelined implementation is slightly better than Beehive, due to the overhead of constructing and deconstructing NoC messages. However, this difference decreases as payload sizes increase since the extra header flits are amortized over a larger payload. The optimized CPU stack remains far below maximum bandwidth even with jumbo frames. The performance difference is especially pronounced at small packet sizes where Beehive is able to sustain echoing 9 Gbps of 64-byte packets (18392 KReq/s) whereas single core Demikernel provides 0.3 Gbps (584 KReq/s), a 31$\times$ speedup.
@@
Latency: For our latency experiment, we use Beehive and a single client thread to ping-pong a single 1-byte UDP packet. We record the latency by tagging the packet with a timestamp when it enters the network stack at the Ethernet parsing layer, taking another timestamp when it finally exits the Ethernet layer on transmit, and recording both timestamps into a log which we read back over the network. The latency through Beehive is 368 ns (92 cycles). Similarly, CALM UDP latency is 362 ns, although their system is less flexible than Beehive.
Figure 9: Packet size vs. goodput for Beehive and Linux TCP send. The (CPU send/FPGA receive) is omitted, as it is approximately the same as (CPU send/CPU receive) due to the CPU send path being the bottleneck.
@@
TABLE III: Energy consumption and goodput for Reed-Solomon encoding using Beehive versus CPU for 1, 2, 3 and 4 application instances.
@@
Energy efficiency 22$\times$ 20$\times$ 20$\times$ 16$\times$
@@
Speedup 7.5$\times$ 7.8$\times$ 7.5$\times$ 7.8$\times$
VII-D TCP throughput
To characterize the throughput performance of our TCP engine, we run a single-connection experiment and measure unidirectional send and receive performance across a range of packet payload sizes. Because Demikernel’s TCP implementation is optimized for latency, it performs worse than Linux on this experiment, so we configure Demikernel to use Linux TCP as its backend. The sending application sits in a tight loop, submitting data into the network stack as fast as possible; the receiver pulls data out of the network stack without doing further processing on it.
@@ VII-D TCP throughput
VII-E Reed-Solomon encoding acceleration
To evaluate Beehive’s scaling architecture, we evaluate a duplicated Reed-Solomon (RS) encoding accelerator on Beehive versus a CPU implementation of the same algorithm in Table LABEL:tab:rs_encode_scaling. The client sends blocks of 4 KB to the encoder using UDP; the accelerator replies with 1K of erasure data. This could be organized into an (8,2) stripe for double fault tolerance. We measured that one instance of the Reed-Solomon encoder can consume data at 15 Gbps; our FPGA has room for four encoder instances, which consume data at 62 Gbps as shown in Table LABEL:tab:rs_encode_scaling. For comparison, we use the open-source Reed-Solomon encoding implementation from BackBlaze [5] running on CPUs which we then duplicate across cores.
@@ VII-E Reed-Solomon encoding acceleration
VII-F Viewstamped replication witness acceleration
@@
Figure 10: Experimental setups for VR evaluation
We next turn to a latency-sensitive application, evaluating Beehive hosting a viewstamped replication (VR) witness appliance. We first evaluate the witness on a single shard. We then take advantage of Beehive’s ability to duplicate both internal components and applications to host a 4-shard witness appliance. We also duplicate protocol tiles to prevent them from becoming a bottleneck.
Setup: For all experiments, we evaluate a three-node VR configuration as shown in Figure 10, with either the FPGA or CPU serving as a witness. Other nodes are run on CPUs. The CPU VR replicas run on Intel Xeon Gold 5218 CPUs. Client threads run on Intel Xeon Gold 6226R CPUs and are closed loop, i.e., only have one outstanding request at a time. The shard leaders are distributed evenly between two CPU machines. Each shard may handle more than one request at a time. The CPU witness(es) run on a separate server to allow us to measure the energy used by a CPU witness appliance. We use UDP as our transport protocol, because VR does not assume reliable message delivery.
Workload and Metrics: We evaluate our VR accelerator with a replicated key-value store application with 64-byte keys and 64-byte values. The workload uses a read-write mix of 90% reads and 10% writes and a uniform key distribution. Input load is increased by increasing the number of clients. Latency is measured at the clients as the time between the initial request and the eventual response. Peak throughput numbers are chosen at the points before the latency begins to spike, an indication that the system is overloading and queues in the system are growing. These points correspond to operational setups where increased latency might be considered acceptable in exchange for better throughput [52, 15, 55].
Results: We plot latency versus throughput for differing numbers of shards in Figure 11. We increase offered load by increasing the number of client threads sending requests to the leader. The results are shown in Figure 11. The system using the FPGA witness can provide up to 1.14$\times$ more per-core throughput and up to 1.13$\times$ lower median latency.
Figure 11: Latency vs. throughput for the VR key-value store workload varying the number of shards and client threads. The FPGA witness consistently outperforms the equivalent CPU cores in both latency and throughput.
TABLE IV: Energy per operation (measured at the witness) and performance metrics (measured at the clients) at the circled points in Figure 11.
@@
For each shard, we take the median energy measurement, throughput, median latency, and 99th-percentile (p99) latency at each circled point in Figure 11. These results are shown in Table IV. The FPGA is between 2.07$\times$ and 2.32$\times$ more energy efficient per operation compared to the CPU while providing better overall throughput and latency to key-value store clients.
VII-G Hardware resource utilization
@@
TABLE V: FPGA resource utilization of selected modules in Beehive and Limago.
LUTs (# / % total) BRAM (# / % total)
Beehive UDP full 58540 / 4.95 41 / 1.90
UDP RX Tile 10054 / 0.85 9.5 / 0.44
       Router 5946 / 0.50 0 / 0
       NoC Message Parsing 897 / 0.07 0 / 0
       UDP RX Processing 2912 / 0.25 9.5 / 0.44
UDP TX Tile 10128 / 0.86 9.5 / 0.44
       Router 5955 / 0.50 0 / 0
       NoC Message Parsing 658 / 0.06 0 / 0
       UDP TX Processing 3105 / 0.26 9.5 / 0.44
Beehive TCP/UDP stack 144491 / 12 84.5 / 4
Beehive TCP Layer 41677 / 3.5 25 / 1.1
TCP RX Processing 10304 / 0.87 9 / 0.4
TCP RX Router 8847 / 0.74 0 / 0
Limago TCP/UDP stack 116948 / 9.9 155 / 7.2
Limago TCP Layer 52134 / 4.4 99 / 4.6
@@
The hardware utilization of the Beehive infrastructure is shown in Table V. For the UDP stack used in Section VII-C, Beehive components use 4% of the LUTs available on the Alveo U200 and 2% of the BRAMs. In a tile, a router uses around 6000 LUTs, twice the size of the UDP processing. For comparison with a more complex module, we include the utilization of the TCP receive path.
We also compare our resource utilization to that of Limago. We find that our design is larger in terms of LUT usage, but smaller in terms of BRAM usage. Most of our usage comes from the routers rather than the protocol logic, indicating that there is a cost to our increased flexibility. However, in the context of total resources available on the FPGA, the extra logic cost is relatively small.
VII-H Flexibility
As a quantitative proxy for flexibility, we count the lines of code (LoC) required to insert an additional instance of an implemented service (network function or application) into the design for our three designs. Results are shown in Table VI.
TABLE VI: Lines of code per new tile instantiation in Beehive for end-to-end applications. XML configuration numbers are given as LoC for declaring the tile plus the LoC to add it as a destination.
Lines of Code
XML Config. Verilog Top Level
Reed-Solomon 25 + 6 13
Viewstamped Replication 18 + (6$\times$# of UDP tiles) 17
VII-I Scalability
We did two experiments to evaluate the scalability of Beehive: one bandwidth-oriented and one hardware resource oriented. For the bandwidth-oriented experiment we repeated the UDP echo experiment in cycle-accurate simulation, duplicating the UDP stack and adding a simple load-balancing tile at the front that splits flows evenly between the stacks. The maximum goodput the load balancer can achieve is 32Gbps for 64-byte UDP packet since each takes 4 cycles to process at the load balancer: 3 for the NoC message and 1 recovery cycle. We hit the maximum possible goodput of the load-balancer of  32Gbps for 64-byte packets. With two stacks, at small packet sizes, we also roughly double the bandwidth as with one stack. This performance difference decreases at larger payload sizes and both stacks converge to the maximum possible goodput of the network link.